OpenStudy (anonymous):

PLEASE HELP A pizza pan is removed at 9:00pm from an oven whose temperature is fixed at 450 F into a room that is constant 72 F. After 5 mins the pizza pan is 300 F (a) determine the time is the temperature of the pan 125 F? (b) Determine the time that needs to elapse before the pan is 230 F (c) what do you notice about the temp. as time passes?

thomaster (thomaster):

Don't know for sure if the temperature decrease would be linear or exponential decay. If it's linear then you can calculate the temperature drop per minute. t0=450F t5=300F So 150F in 5 minutes 150/5 is 30F drop per minute Formula = 450 - 30t t is the minutes So to know t for temp = 125F: 450-30t = 125 450 - 125 = 30t --> 325 = 30t t= 325/30= 10.8 minutes Same thing for B C: the temperature goes down BUT: i don't think that the temperature would go down in a linear way. I think it will go fast in the beginning and slower and slower after that. So exponential decay but im way to lazy to calculate that shyt its 3am >.<

OpenStudy (anonymous):

it is not linear

OpenStudy (anonymous):

first off work with the temperature differences, that is \(450-72=378\) and \(300-72=228\) so in 5 minutes the temperature difference goes from 378 to 228

OpenStudy (anonymous):

your job is to model exponential decay from \((0,378)\) to \((5,228)\) the ratio is \(\frac{228}{378}=\frac{38}{63}\) the simple way to do this is to write the equation for the temperature differences is \[378\times\left( \frac{38}{63}\right)^{\frac{t}{5}}\] although you can also model it as \[378\times e^{kt}\] but in that case you have to find \(k\)

OpenStudy (anonymous):

if you want \(k\) you can put \[228=378e^{5k}\] making \[e^{5k}=\frac{38}{63}\] and so \[k=\frac{\ln(\frac{38}{63})}{5}\]

OpenStudy (anonymous):

so how would it look all together put? would it look like 378 * e^(ln(38/63))/5)

OpenStudy (anonymous):

yeah but you sort of want a number in the exponent that is, compute \[k=-.10111\] and your equation for the differences is \[378e^{-.10111t}\]

OpenStudy (anonymous):

then it will be 125 degrees when the temperature difference is \[125-72=53\] solve \[378e^{-.10111t}=53\] for \(t\)

OpenStudy (anonymous):

how will you get get t then? do you have to do the natural log?

OpenStudy (anonymous):

yes, divide by \(378\) then take the log, then divide by \(-.10111\)

OpenStudy (anonymous):

\[378e^{-.10111t}-53\] \[e^{-.10111t}=\frac{53}{378}\] \[-.10111t=\ln(\frac{53}{378})\] \[t=\ln(\frac{53}{378})\div-.10111\]

OpenStudy (anonymous):

I got 19.43

OpenStudy (anonymous):

how will this get the time? for part a?

OpenStudy (anonymous):

that is part A

OpenStudy (anonymous):

yeah I just realized it!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so for part b the starting equation will be 230=72+450e^-0.10111 t and so I will have to find t

OpenStudy (anonymous):

no \[230=72+378\times e^{-0.10111 t}\]

OpenStudy (anonymous):

I got 8.627

OpenStudy (anonymous):

would the temp of the pan be approaching the room tem. 72 F